OpenStudy (anonymous):

Which of the following is the slope between the two points (1, 5) and (6, 4)? a.1/5 b.- 1/5 c. 5 d.-5

OpenStudy (anonymous):

Recall the formula for the slope between two points: \[m= \Delta y / \Delta x=(y_1 - y_2)/(x_1 - x_2)\]Now you can plug in your points to find the slope.

OpenStudy (texaschic101):

do you understand what vinny is saying ?

OpenStudy (texaschic101):

slope(m) = (y2 - y1) / (x2 - x1) (1,5) ... x1 = 1, y1 = 5 (6,4)...x2 = 6, y2 = 4 Now just sub those into the formula and you will find your slope